OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -2x - y < -6. A. Dashed line with shading on the side that includes the origin. B. Solid line with shading on the side that does not include the origin. C. Dashed line with shading on the side that does not include the origin. D. Solid line with shading on the side that includes the origin. I know it would be a dashed line but the point of origin kind of through's me off, can someone help me?? Please!! :)

OpenStudy (anonymous):

Connie needs to burn at least 500 calories at the gym today. She can burn 20 calories per minute on the treadmill (t) and 10 calories per minute on the elliptical trainer (e). Which inequality below can be used to find the possible combinations of time she could spend on each machine and still meet her goal of burning at least 500 calories? A. 20t - 10e >= 500 B. 20t + 10e <= 500 C. 20t - 10e <= 500 D. 20t + 10e >= 500 I think this one is B?

OpenStudy (anonymous):

plug in x = 0, y = 0 so -2x - y < -6 becomes 0 - 0 < -6 or 0 < -6 is this true? ................. no ....... so the shading does not include (0,0) and the answer is C

OpenStudy (anonymous):

ohh okay, can you help with the other question to?

OpenStudy (anonymous):

Actually, could you check my answers for a few more questions? Ill post the on yahoo answers to make it easier if your willing to do that for me? let me know! :)

OpenStudy (anonymous):

Here is the link the yhaoo answers: http://answers.yahoo.com/question/index?qid=20120118081805AAplf2p

OpenStudy (anonymous):

Please check it out and see if im correct

OpenStudy (anonymous):

I would truly and great-fully appreciate it